QUESTION 6

  1. What factor(s), besides the NPV of an investment, should managers consider in a capital budgeting decision?

    The effect of the investment on the manager's short-term incentives

    The effect of the investment on the company's reputation

    Tax shields related to depreciation that would result from the investment

    All three of the other answers

2 points   

QUESTION 7

  1. Which of the following three investments will have the highest NPV? All three cases require an initial investment of $45,000, and the required rate of return is 10%.

    A B C
    Year 1 net cash in 12,000 27,000 42,000
    Year 2 net cash in 22,000 27,000 32,000
    Year 3 net cash in 32,000 27,000 22,000
    Year 4 net cash in 42,000 27,000 12,000
    Total cash flow    108,000 108,000 108,000

    Investment B

    Investment C

    All three investments will have the same NPV

    Investment A

2 points   

QUESTION 8

  1. Sam's Manufacturing Company currently makes 100 units of a necessary component. Management is considering outsourcing this component for a cost of $1,700 per unit. Sam's incurs the following total production costs:

    Direct Materials $80,000
    Direct Labor 23,000
    Variable Overhead 37,000
    Fixed Overhead 24,000


    If production is outsourced, none of the fixed overhead costs will be eliminated. How would profits be impacted if Sam's bought the component?

    Profits would go up by $67,000

    Profits would go down by $6,000

    Profits would go down by $30,000

    Profits would go up by $6,000

2 points   

QUESTION 9

  1. Adler Company manufactures 10,000 units of a product for a unit sales price of $88. Variable costs are $50 per unit and annual fixed manufacturing costs total $240,000. The company has a one-time opportunity to sell an additional 3,000 units at $70 each in a foreign market. If the company has sufficient capacity to produce the additional units, acceptance of the special order would affect income as follows:

    Income would decrease by $12,000.

    Income would increase by $12,000.

    Income would increase by $210,000.

    Income would increase by $60,000.

2 points   

QUESTION 10

  1. Globiane, Inc. has designed a new product that could be sold for $1,000. If management requires a profit of 30% of the selling price, what is the highest cost (target cost) management would accept to make this product?

    $700

    $300

    $1,429

    $3,333
Answer #1

6.All three of the other answer.

7.Answer is INVESTMENT C

YEAR A B C
Year 1 net cash in 12,000.00 0.909 10,908 27,000.00 0.909 24,543 42,000.00 0.909 38,178
Year 2 net cash in 22,000.00 0.826 18,172 27,000.00 0.826 22,302 32,000.00 0.826 26,432
Year 3 net cash in 32,000.00 0.751 24,032 27,000.00 0.751 20,277 22,000.00 0.751 16,522
Year 4 net cash in 42,000.00 0.683 28,686 27,000.00 0.683 18,441 12,000.00 0.683 8,196
Present Value of Inflow 81,798 85,563 89,328
Present value of Outflow 45000 45000 45000
Net Present Value 36,798 40,563 44,328

8.)

Direct Materials    80,000.00
Direct Labor    23,000.00
Variable Overhead    37,000.00
Total Costs    140,000.00
Cost Per unit of Make        1,400.00

Profit would go down by = (1,700 - 1,400) x 100 = 30,000

9)Income would increase by = (70 - 50) x 3,000

=60,000

10)Profit = 1,000 x 30% = 300

So Target Cost can be 1,000 - 300 = 700